Suppose the following system reaches equilibrium. N2(g) + O2(g) ←→ 2NO(g) Analysis of the mixture in a 1-L flask gives the following results: N2 = 0.50 mol, O2 = 0.50 mol, and NO = 0.020 mol. Calculate Keq for the reaction.
1
Expert's answer
2020-04-28T12:19:47-0400

Amount of "N_2 = 0.50\ mol, O_2 = 0.50\ mol," and "NO=0.020\ mol" in a "1" litre flask.

And,Concentration is amount of solute in "1" litre of solution.

Therefore,"[N_2]=0.5\ mol/l;[O_2]=0.5\ mol/l" ";[NO]=0.020 \ mol/l"

For the given reaction,"N_2(g) + O_2(g) \u2190\u2192 2NO(g)"

"K_{eq}=\frac{[NO]^2}{[N_2][O_2]}=\frac{(0.020)^2}{0.5\times0.5}=0.080"